The equilibrium constant, Kc, for the following reaction is 1.80×10-4 at 298 K.

NH4HS(s) H3(g) +H2S(g) Calculate the equilibrium concentration of H2S when 0.261 moles of NH4HS(s) are introduced into a 1.00 L vessel at 298 K. [H2S] =
Answer #1

since volume of the container is 1L moles = molarity   

construct the ICE table

NH4HS (s) <===> NH3 (g) + H2S (g)

I 0.261 0 0

C -x +x +x

E 0.261-x +x +x

now write the equilibrium expression

while writing the equilibrium expression we will ignore the solide part which is NH4HS so

Kc = [NH3] [H2S]

1.80×10-4 = [x] [x]

x2 = 1.80×10-4

x = \sqrt{}1.80×10-4

x = 0.0134 M

[H2S] = 0.0134 M